# SqueezeMe: Mobile-Ready Distillation of Gaussian Full-Body Avatars

### v3

@@ Abstract.
Gaussian-based human avatars have achieved an unprecedented level of visual fidelity. However, existing approaches based on high-capacity neural networks typically require a desktop GPU to achieve real-time performance for a single avatar, and it remains non-trivial to animate and render such avatars on mobile devices including a standalone VR headset due to substantially limited memory and computational bandwidth. In this paper, we present SqueezeMe, a simple and highly effective framework to convert high-fidelity 3D Gaussian full-body avatars into a lightweight representation that supports both animation and rendering with mobile-grade compute. Our key observation is that the decoding of pose-dependent Gaussian attributes from a neural network creates non-negligible memory and computational overhead. Inspired by blendshapes and linear pose correctives widely used in Computer Graphics, we address this by distilling the pose correctives learned with neural networks into linear layers. Moreover, we further reduce the parameters by sharing the correctives among nearby Gaussians. Combining them with a custom splatting pipeline based on Vulkan, we achieve, for the first time, simultaneous animation and rendering of 3 Gaussian avatars in real-time (72 FPS) on a Meta Quest 3 VR headset.
@@
ccs: Computing methodologies Reconstructionccs: Computing methodologies Animation
@@ 1. Introduction
Recent advances in accelerating the rendering of photorealistic 3D scenes based on NeRF (Chen et al., 2023a; Reiser et al., 2023, 2024) and 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023; Niedermayr et al., 2024) have shown promising results. Moreover, real-time rendering of volumetric videos on mobile devices has become possible (Wang et al., 2024b). However, these approaches are not directly applicable to avatars, as they require support for not only static scenes or the playback of pre-recorded sequences, but also novel animations computed on-the-fly for real-time driving of avatars (Bagautdinov et al., 2021). Existing approaches that can achieve photorealism rely on high-capacity neural networks to decode non-rigid correctives given a driving signal such as body pose parameters (Li et al., 2024; Wang et al., 2022; Svitov et al., 2024). Running such a decoder for every frame creates non-negligible memory and computational overhead, hindering the deployment of photorealistic avatars on device.
@@ 2.1. Photorealistic Full-Body Avatars
Creating photrealistic avatars has been a long-standing problem in computer graphics and vision. Earlier works attempt modeling clothed humans by deforming a single template mesh such as SMPL (Loper et al., 2015a) with texture maps (Alldieck et al., 2018b, a; Bagautdinov et al., 2021). While the mesh-based approaches work well for tight clothing and skin regions, hair is typically difficult to handle due to topological constraints and difficulty in handling opacity. Avatars based on Signed Distance Fields (SDF) (Saito et al., 2021; Wang et al., 2022; Guo et al., 2023; Chen et al., 2024) can handle topology changes, but have difficulty handling intricate geometric details such as hair. Volumetric representations such as Neural Volumes (Lombardi et al., 2019) or NeRF (Mildenhall et al., 2021) address the limitation of meshes by explicitly integrating opacity in volumetric rendering equations. These approaches are later expended to human modeling (Weng et al., 2020; Peng et al., 2021b, a; Liu et al., 2021; Su et al., 2021; Jiang et al., 2022; Li et al., 2022; Weng et al., 2022). While these volumetric representations are typically slow to render due to expensive evaluation, follow-up works accelerate the rendering for real-time inference by utilizing primitive approximation of volumetric fields with voxels (Lombardi et al., 2021; Remelli et al., 2022; Chen et al., 2023b) or point clouds (Zheng et al., 2023; Su et al., 2023; Prokudin et al., 2023). However, these approaches still struggle with modeling detailed shape and appearance. More recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) achieves better trade-off for rendering speed and fidelity by approximating radiance fields with anisotropic Gaussians. Full-body avatars based on 3DGS achieve state-of-the-art performance (Hu and Liu, 2024; Zielonka et al., 2023; Li et al., 2024; Qian et al., 2024; Pang et al., 2024). While these approaches based on 3DGS achieves real-time animation and rendering on desktop-grade GPUs, decoding pose correctives per frame remains prohibitively slow or intractable to run on mobile compute due to high memory and computation requirements. In contrast, our work proposes the distillation of these high-capacity CNN decoders into light weight linear layers with significant memory reduction by corrective sharing, enabling the animation and rendering of pose-dependent animatable Gaussian avatars.
@@ 2.2. Efficient Volumetric Rendering
Efficient rendering of radiance fields are also extensively studied. Earlier works utilize multiplane image (MPI)  (Wizadwongsa et al., 2021) or sparse voxel grids (Hedman et al., 2021) to drastically reduce the number of evaluations per ray. More recent works further reduces the evaluation time to once per ray by extracting surface mesh (Chen et al., 2023a; Reiser et al., 2023; Yariv et al., 2023; Duckworth et al., 2024; Guédon and Lepetit, 2024). 3DGS rendering can also be accelerated by integer quantization, vector quantization, and hash grids to reduce the size footprint of 3DGS as demonstrated in (Fan et al., 2023; Niedermayr et al., 2024; Lee et al., 2024). These approaches are only applicable to static scenes, and more recent works support efficient rendering of 4D dynamic scenes using triplane (Wu et al., 2024), voxels (Wang et al., 2024a), depth peeling (Xu et al., 2024), and dynamic 3D Gaussians (Sun et al., 2024). However, they are either non-trivial to stream and render on mobile devices (Xu et al., 2024; Sun et al., 2024; Wu et al., 2024) or prone to blurry results (Wang et al., 2024a). V3 (Wang et al., 2024b) achieves mobile streaming of dynamic Gaussian splatting by converting dynamic Gaussians to 2D image sequences and applying video codec. Unlike these methods, which support only playback of pre-recorded sequences, our approach compresses the basis of correctives and effectively supports novel animations of photorealistic avatars with on-the-fly driving signals with mobile-grade compute.
@@ 2.3. Efficient Animatable Avatars
Improving the efficiency of rendering and animation of photorealistic avatars is an active research topic in the community. Pixel Codec Avatars (Ma et al., 2021) presents a method to reduce the memory and computation overhead of head avatars by offloading view-dependent texture decoding to per-pixel computation via neural deferred rendering (Thies et al., 2019) with low-resolution latent codes and high-resolution static maps. MoRF utilizes neural deferred shading for full-body avatar animation and rendering (Bashirov et al., 2024). However, due to the limited compute on mobile, the framerate and resolution are limited to 30 FPS with $640\times 640$ on Qualcomm Snapdragon. Unfortunately, this is insufficient for VR, where we need at least 72 FPS with $2K$ resolution for two eyes. SplattingAvatar (Shao et al., 2024) is a hybrid representation of Gaussian Splats embedded on a tracked mesh, achieving 30 FPS for a single avatar on iPhone 13. However, this approach does not account for pose-dependent correctives, limiting its animation fidelity. The closest to our work are Gaussian Blendshapes (Ma et al., 2024) and Gaussian Eigen Models (Zielonka et al., 2024), where the linear basis of Gaussian parameters in head avatars is extracted from pretrained CNN decoders. However, we find that simply extracting linear basis is still not sufficient to meet memory and computational requirements on mobile. Moreover, unlike head only avatars, full-body avatars need to explicitly account for pose-dependent deformations during the distillation process. In contrast, we present, for the first time, a complete system to enable real-time animation and rendering of full-body Gaussian avatars that meet requirements for VR rendering.
@@ 3.1. Gaussian Splatting
where $N$ is the number of Gaussians visible to pixel $p$, and the view-dependent color $c_{i}$ is computed using spherical harmonics.
@@ 4.1. Compact 3D Gaussian Avatar
Our loss function is a weighted-sum of several terms: $\mathcal{L}_{photo}$ and $\mathcal{L}_{lpips}$ are the photometric (L1) and LPIPS (Zhang et al., 2018) distance between the ground-truth and predicted image. $\mathcal{L}_{kpt}$ is the L1 loss, and the body pose keypoints. To reduce ”runaway” Gaussians that drift far from their initial position, $\mathcal{L}_{offset}$ is an L1 loss that is maximized when the mean $\mu$ of each Gaussian in $G$ is unchanged from $\mu$ where the Gaussian was initialized at the start of training. To reduce unwanted holes in the avatar, we introduce $\mathcal{L}_{\alpha}$, which is an L1 loss that is maximized when the $\alpha$-transparency map generated by the model matches the Sapiens (Khirodkar et al., 2024) segmentation mask; the intuition is that the avatar should be opaque and background pixels should be transparent. When rendering Gaussians on a mobile, it is cheaper if each Gaussian is visible from fewer pixels (Niemeyer et al., 2024). To reduce the number of pixels to which Gaussian is applied during rendering, we adopt the loss $\mathcal{L}_{\sigma}$, which uses L1 to minimize the average size of the Gaussians, and we adopt $\mathcal{L}_{opacity}$, which uses L1 to minimize the average opacity of the Gaussians. The weights for the sum are set to $\lambda_{photo}=1$, $\lambda_{lpips}=0.1$, $\lambda_{opacity}=0.01$, $\lambda_{scale}=1$, $\lambda_{offset}=1$, $\lambda_{alpha}=0.1$, and $\lambda_{keypoint}=0.1$. The loss is computed as
@@ 4.2. Mobile-ready Distillation
For one frame, the linearized decoder’s input is a pose vector $p$, which the concatenation of the face encoder’s embedding and the body pose in quaternion form. We use PCA to compress the input vector $\theta$ to a total of 64d, which we then augment with a columns of ones to form a design matrix $C$. We compute the linear layer in the least-squares sense using the normal equations i.e., we form $(C^{T}C)^{-1}C^{T}$ as the pseudo-inverse of the design matrix $C$ which is then multiplied by the decoder’s output. In order to reduce the decoder output space, as additional optimization, we also use PCA to compress the spherical harmonics from 27 values to 6 values, which yields a 60381x16 tensor. We describe the linear distillation more precisely in Algorithm 1.
@@
The distilled decoder is a 2-layer model. The first layer is a linear layer that takes a 64d compressed vector and outputs a 60381x16 vector, and the layer has 64x60381x16 parameters. The second layer takes 6 of the 16 values from the first layer and expands them into 27 spherical harmonics, and the layer has 6x27 parameters. By far, the dominant cost is the first layer, which we find takes 5 ms to run (with quantization) on the neural processing unit (NPU) of the XR2G2 chip in the Meta Quest 3 VR headset.
@@
While the decoder from Equation 4 produces a size 256x256x37 output, we modify the decoder to produce a smaller 64x64x37 output. For notation, we call the modified decoder $D_{GCS}$, where GCS is short for Gaussian corrective sharing. This reduces the number of correctives from 65536 to 4096. During training, we take the output of $D_{GCS}$ and use Nearest interpolation to upscale it from 64x64 to 256x256 = 65536 correctives. This has the effect of ”sharing” correctives for each 4x4 neighborhood of Gaussians in UV-space. For notation, we abbreviate Nearest upscaling as $Up$. During training, we compute the Gaussians for one frame of animation as
@@ 4.3. Rendering
We have implemented our visualization system in Vulkan to take advantage of the hardware rasterization for accelerated Gaussian splatting, as in (Xu and Yu, 2024; Niedermayr et al., 2024). This is crucial for achieving high performance in compute-constrained setups like mobile devices. We now provide a brief overview of the rendering pipeline in VR. We first run a compute step that animates the avatars with linear blend skinning as proposed in Fig. 2. Next, we perform per-primitive culling to remove Gaussians that are outside the field-of-view for both stereo views in VR. Then, we project and sort the remaining Gaussians according to their view depth and reuse the sorted indices across both eye views. Finally, we expand the projected Gaussians into quad primitives with corresponding colors and opacities, which are rasterized and blended via a traditional graphics pipeline. All of the stages are implemented in standard Vulkan compute shaders except for the last (rasterization) stage, which uses a combination of a vertex and a fragment shader.
@@ 5. Experimental Setup
Table 1. Main results. We evaluate the impact of linearization and number of correctives on quality. Results are at the original image resolution, with images cropped to rectangles based on segmentation of the avatar. Results are averaged over 4 identities. Latency is for the decoder only.
Model # Gaussians # Correctives Linearized L1 $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
@@
SqueezeMe (initial) 60k 60k 0.037 0.145 24.728 0.625
SqueezeMe (GCS) 60k 4k 0.037 0.146 24.763 0.623
SqueezeMe (Linearized) 60k 60k $\checkmark$ 0.040 0.150 24.138 0.618 5.0 ms
SqueezeMe (GCS + Linearized) 60k 4k $\checkmark$ 0.040 0.150 24.115 0.615 0.45 ms
SqueezeMe (no decoder) 60k 0 0.041 0.164 23.626 0.607 0
@@
In all of our evaluations, we are performing novel-view synthesis, i.e. we are using camera positions that were not in training set. All results are averaged across 4 identities and 5 cameras per identity. The test set is held-out from the dataset that is used for training, linear distillation, and quantization.
For on-device demo, we use a Meta Quest 3 VR headset. The Quest 3 is equipped with a Snapdragon XR2 Gen 2 chip, which integrates a CPU, a GPU, and a neural processing unit (NPU) (Qualcomm, ). Unlike desktop GPUs, Snapdragon does not support CUDA; however, its mobile GPU is programmable via the Vulkan API, and its NPU can execute quantized neural networks exported from frameworks like PyTorch.

### v4

@@ Abstract.
Gaussian-based human avatars have achieved an unprecedented level of visual fidelity. However, existing approaches based on high-capacity neural networks typically require a desktop GPU to achieve real-time performance for a single avatar, and it remains non-trivial to animate and render such avatars on mobile devices including a standalone VR headset due to substantially limited memory and computational bandwidth. In this paper, we present SqueezeMe, a simple and highly effective framework to convert high-fidelity 3D Gaussian full-body avatars into a lightweight representation that supports both animation and rendering with mobile-grade compute. Our key observation is that the decoding of pose-dependent Gaussian attributes from a neural network creates non-negligible memory and computational overhead. Inspired by blendshapes and linear pose correctives widely used in Computer Graphics, we address this by distilling the pose correctives learned with neural networks into linear layers. Moreover, we further reduce the parameters by sharing the correctives among nearby Gaussians. Combining them with a custom splatting pipeline based on Vulkan, we achieve, for the first time, simultaneous animation and rendering of 3 Gaussian avatars in real-time (72 FPS) on a Meta Quest 3 VR headset. Demo videos are available at https://forresti.github.io/squeezeme.
@@
submissionid: 192copyright: ccconference: ; August 10–14, 2025; Vancouver, BCjournal: TOGccs: Computing methodologies Reconstructionccs: Computing methodologies Animation
@@ 1. Introduction
Recent advances in accelerating the rendering of photorealistic 3D scenes based on NeRF (Chen et al., 2023a; Reiser et al., 2023, 2024) and 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023; Niedermayr et al., 2024) have shown promising results. Moreover, real-time rendering of volumetric videos on mobile devices has become possible (Wang et al., 2024b). However, these approaches are not directly applicable to avatars, because avatars require support for not only static scenes or the playback of pre-recorded sequences, but also novel animations computed on-the-fly for real-time driving of avatars (Bagautdinov et al., 2021). Existing approaches that can achieve photorealism rely on high-capacity neural networks to decode non-rigid correctives given a driving signal such as body pose parameters (Li et al., 2024; Wang et al., 2022; Svitov et al., 2024). Running such a decoder for every frame creates non-negligible memory and computational overhead, hindering the deployment of photorealistic avatars on device.
@@ 2.1. Photorealistic Full-Body Avatars
Creating photrealistic avatars has been a long-standing problem in computer graphics and vision. Earlier works attempt modeling clothed humans by deforming a single template mesh such as SMPL (Loper et al., 2015a) with texture maps (Alldieck et al., 2018b, a; Bagautdinov et al., 2021). While the mesh-based approaches work well for tight clothing and skin regions, hair is typically difficult to handle due to topological constraints and difficulty in handling opacity. Avatars based on Signed Distance Fields (SDF) (Saito et al., 2021; Wang et al., 2022; Guo et al., 2023; Chen et al., 2024; Jiang et al., 2022a; Zhu et al., 2024) can handle topology changes, but have difficulty handling intricate geometric details such as hair. Volumetric representations such as Neural Volumes (Lombardi et al., 2019) or NeRF (Mildenhall et al., 2021) address the limitation of meshes by explicitly integrating opacity in volumetric rendering equations. These approaches are later expended to human modeling (Weng et al., 2020; Peng et al., 2021b, a; Liu et al., 2021; Su et al., 2021; Jiang et al., 2022b; Li et al., 2022; Weng et al., 2022). While these volumetric representations are typically slow to render due to expensive evaluation, follow-up works accelerate the rendering for real-time inference by utilizing primitive approximation of volumetric fields with voxels (Lombardi et al., 2021; Remelli et al., 2022; Chen et al., 2023b) or point clouds (Zheng et al., 2023a; Su et al., 2023; Prokudin et al., 2023). However, these approaches still struggle with modeling detailed shape and appearance. More recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) achieves better trade-off for rendering speed and fidelity by approximating radiance fields with anisotropic Gaussians. Full-body avatars based on 3DGS achieve state-of-the-art performance (Hu and Liu, 2024; Zielonka et al., 2023; Li et al., 2024; Qian et al., 2024; Pang et al., 2024; Hu et al., 2024). While these approaches based on 3DGS achieves real-time animation and rendering on desktop-grade GPUs, decoding pose correctives per frame remains prohibitively slow or intractable to run on mobile compute due to high memory and computation requirements. In contrast, our work proposes the distillation of these high-capacity CNN decoders into lightweight linear layers with significant memory reduction by corrective sharing, enabling the animation and rendering of pose-dependent animatable Gaussian avatars.
@@ 2.2. Efficient Volumetric Rendering
Efficient rendering of radiance fields are also extensively studied. Earlier works utilize multiplane image (MPI)  (Wizadwongsa et al., 2021) or sparse voxel grids (Hedman et al., 2021) to drastically reduce the number of evaluations per ray. More recent works further reduce the evaluation time to once per ray by extracting the surface mesh (Chen et al., 2023a; Reiser et al., 2023; Yariv et al., 2023; Duckworth et al., 2024; Guédon and Lepetit, 2024). 3DGS rendering can also be accelerated by integer quantization, vector quantization, and hash grids to reduce the size footprint of 3DGS as demonstrated in (Fan et al., 2023; Niedermayr et al., 2024; Lee et al., 2024). These approaches are only applicable to static scenes, and more recent works support efficient rendering of 4D dynamic scenes using triplane (Wu et al., 2024), voxels (Wang et al., 2024a; Lin et al., 2023), depth peeling (Xu et al., 2024a), and dynamic 3D Gaussians (Sun et al., 2024; Yang et al., 2024; Xu et al., 2024b). However, they are either non-trivial to stream and render on mobile devices (Xu et al., 2024a; Sun et al., 2024; Wu et al., 2024) or prone to blurry results (Wang et al., 2024a). V3 (Wang et al., 2024b) achieves mobile streaming of dynamic Gaussian splatting by converting dynamic Gaussians to 2D image sequences and applying video codec. Unlike these methods, which support only playback of pre-recorded sequences, our approach compresses the basis of correctives and effectively supports novel animations of photorealistic avatars with on-the-fly driving signals with mobile-grade compute.
@@ 2.3. Efficient Animatable Avatars
Improving the efficiency of rendering and animation of photorealistic avatars is an active research topic in the community. Pixel Codec Avatars (Ma et al., 2021) presents a method to reduce the memory and computation overhead of head avatars by offloading view-dependent texture decoding to per-pixel computation via neural deferred rendering (Thies et al., 2019) with low-resolution latent codes and high-resolution static maps. MoRF utilizes neural deferred shading for full-body avatar animation and rendering (Bashirov et al., 2024). However, due to the limited compute on mobile, the framerate and resolution are limited to 30 FPS with $640\times 640$ on Qualcomm Snapdragon. Unfortunately, this is insufficient for VR, where we need at least 72 FPS with $2K$ resolution for two eyes. SplattingAvatar (Shao et al., 2024) is a hybrid representation of Gaussian Splats embedded on a tracked mesh, achieving 30 FPS for a single avatar on iPhone 13. However, this approach does not account for pose-dependent correctives, limiting its animation fidelity. The closest to our work are Gaussian Blendshapes (Ma et al., 2024; Yan et al., 2025) and Gaussian Eigen Models (Zielonka et al., 2024), where the linear basis of Gaussian parameters in head avatars is extracted from pretrained CNN decoders. However, we find that simply extracting linear basis is still not sufficient to meet memory and computational requirements on mobile. Moreover, unlike head only avatars, full-body avatars need to explicitly account for pose-dependent deformations during the distillation process. In contrast, we present, for the first time, a complete system to enable real-time animation and rendering of full-body Gaussian avatars that meet requirements for VR rendering.
@@ 3.1. Gaussian Splatting
where $N$ is the number of Gaussians visible to pixel $\rho$, and the view-dependent color $c_{i}$ is computed using spherical harmonics.
@@ 4.1. Compact 3D Gaussian Avatar
Our loss function is a weighted-sum of several terms: $\mathcal{L}_{photo}$ and $\mathcal{L}_{lpips}$ are the photometric (L1) and LPIPS (Zhang et al., 2018) distance between the ground-truth and predicted image. $\mathcal{L}_{kpt}$ is the L1 loss, and the body pose keypoints. To reduce ”runaway” Gaussians that drift far from their initial position, $\mathcal{L}_{offset}$ is an L1 loss that is maximized when the mean $\mu$ of each Gaussian in $G$ is unchanged from $\mu$ where the Gaussian was initialized at the start of training. To reduce unwanted holes in the avatar, we introduce $\mathcal{L}_{\alpha}$, which is an L1 loss that is maximized when the $\alpha$-transparency map generated by the model matches the Sapiens (Khirodkar et al., 2024) segmentation mask; the intuition is that the avatar should be opaque and background pixels should be transparent. When rendering Gaussians on a mobile, it is cheaper if each Gaussian is visible from fewer pixels (Niemeyer et al., 2024). To reduce the number of pixels to which Gaussian is applied during rendering, we adopt the loss $\mathcal{L}_{\sigma}$, which uses L1 to minimize the average size of the Gaussians, and we adopt $\mathcal{L}_{opacity}$, which uses L1 to minimize the average opacity of the Gaussians. We visualize the effects of the losses in the supplementary material. The weights for the sum are set to $\lambda_{photo}=1$, $\lambda_{lpips}=0.1$, $\lambda_{opacity}=0.01$, $\lambda_{scale}=1$, $\lambda_{offset}=1$, $\lambda_{alpha}=0.1$, and $\lambda_{kpt}=0.1$. The loss is computed as
@@ 4.2. Mobile-ready Distillation
For one frame, the linearized decoder’s input is a pose vector $p$, which the concatenation of the face encoder’s embedding and the body pose in quaternion form. We use PCA to compress the input vector $\theta$ to a total of 64 dimensions, which we then augment with a columns of ones to form a design matrix $C$. We compute the linear layer in the least-squares sense using the normal equations i.e., we form $(C^{T}C)^{-1}C^{T}$ as the pseudo-inverse of the design matrix $C$ which is then multiplied by the decoder’s output. We describe the linear distillation more precisely in Algorithm 1.
@@
The distilled decoder is a 2-layer model. The first layer is a linear layer that takes a 64d compressed vector and outputs a 60381x16 vector, and the layer has 64x60381x16 parameters. The second layer takes 6 of the 16 values from the first layer and expands them into 27 spherical harmonics, and the layer has 6x27 parameters. The remaining 10 channels are used for the Gaussian geometric parameters such as rotation, scale, and translation. By far, the dominant cost is the first layer, which we find takes 5 ms to run (with quantization) on the neural processing unit (NPU) of the XR2G2 chip in the Meta Quest 3 VR headset.
@@
While the decoder from Equation 4 produces a size 256x256x37 output, we modify the decoder to produce a smaller 64x64x37 output. For notation, we call the modified decoder $D_{GCS}$, where GCS is short for Gaussian corrective sharing. This reduces the number of correctives from 65536 to 4096. During training, we take the output of $D_{GCS}$ and use Nearest interpolation to upscale it from 64x64 to 256x256 = 65536 correctives. [^1] We experimented with both bilinear interpolation and nearest interpolation. See supplementary material for details. This has the effect of ”sharing” correctives for each 4x4 neighborhood of Gaussians in UV-space. For notation, we abbreviate Nearest upscaling as $Up$. During training, we compute the Gaussians for one frame of animation as
@@ 4.3. Rendering
We have implemented our visualization system in Vulkan to take advantage of the hardware rasterization for accelerated Gaussian splatting, as in (Xu and Yu, 2024; Niedermayr et al., 2024). This is crucial for achieving high performance in compute-constrained setups like mobile devices. We now provide a brief overview of the rendering pipeline in VR. We first run a compute step that animates the avatars with linear blend skinning as proposed in Fig. 2. Next, we perform per-primitive culling to remove Gaussians that are outside the field-of-view for both stereo views in VR. Then, we project and sort the remaining Gaussians according to their view depth and reuse the sorted indices across both eye views. Finally, we expand the projected Gaussians into quad primitives with corresponding colors and opacities, which are rasterized and blended via a traditional graphics pipeline. The rationale for this final step is that Gaussian primitives are ellipsoids, but traditional graphics hardware is optimized for triangles and quadrilaterals. All of the stages are implemented in standard Vulkan compute shaders except for the last (rasterization) stage, which uses a combination of a vertex and a fragment shader.
@@ 5. Experimental Setup
Table 1. Main results. We evaluate the impact of linearization and number of correctives on quality. Results are at the original image resolution, with images cropped to rectangles based on segmentation of the avatar. Results are averaged over 4 identities. Latency is for the decoder only. One of the identities in Linear from scratch diverged due to instability and the reported number is the average across the remaining 3 identities.
Model # Gaussians # Correctives L1 $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
@@
SqueezeMe (initial) 60k 60k 0.036 0.146 24.992 0.638
SqueezeMe (GCS) 60k 4k 0.036 0.147 25.024 0.636
SqueezeMe (Linearized) 60k 60k 0.039 0.150 24.411 0.632 5.0 ms
SqueezeMe (GCS + Linearized) 60k 4k 0.039 0.151 24.370 0.629 0.45 ms
SqueezeMe (no decoder) 60k 0 0.040 0.164 23.905 0.622 0
Linear from scratch 60k 60k 0.044* 0.215* 23.454* 0.616* 5.0 ms
@@
Training and evaluation use two separate sets of poses from distinctive animation segments. All results tables, figures, and videos are computed on the evaluation set. Further, in all of our evaluations, we are performing novel-view synthesis, i.e. we are using camera positions that were not in training set. The evaluation set is held-out from the dataset that is used for training, linear distillation, and quantization. For each subject, we use 2000-3000 frames per subject for training and 2000 held-out frames for evaluation. For distillation, we use 512 frames from the training set. All results are averaged across 4 identities and 5 cameras per identity.
For on-device demonstration, we use a Meta Quest 3 VR headset. The Quest 3 is equipped with a Snapdragon XR2 Gen 2 chip, which integrates a CPU, a GPU, and a neural processing unit (NPU) (Qualcomm, ). Unlike desktop GPUs, Snapdragon does not support CUDA; however, its mobile GPU is programmable via the Vulkan API, and its NPU can execute quantized neural networks exported from frameworks like PyTorch.